OpenStudy (anonymous):

A train is traveling at 100km/h during a fog. as the train comes out into a clear stretch, the engineer suddenly sees a car stalled on the tracks 0.50km away. she applies the brakes which are capable of giving the train an acceleration of -1.0m/s^2. a how long a time will it take the train to stop?

OpenStudy (anonymous):

for the time I got 31.6s, is that right?

OpenStudy (theeric):

I used \(a=\dfrac{v_f-v_i}{\Delta t}\) \(\implies \Delta t = \dfrac{v_f-v_i}{a}= \dfrac{-v_i}{a}\)

OpenStudy (theeric):

A more fun problem is "will the train hit the car?"

OpenStudy (anonymous):

Thank you so much!

OpenStudy (theeric):

You're welcome! :) Let me know if you have any more questions on this!